### v2

# Pruning Large Language Models with Semi-Structural Adaptive Sparse Training

Abstract
The tremendous success of Large Language Models (LLMs) across various complex tasks relies heavily on their substantial scale, which raises challenges during model deployment due to their large memory consumption. Recently, numerous studies have attempted to compress LLMs using one-shot pruning methods. However, these methods often experience considerable performance degradation on complex language understanding tasks, calling into question the feasibility of pruning in LLMs. To address this issue, we propose a pruning pipeline for semi-structured sparse models via retraining, termed Adaptive Sparse Trainer (AST). Unlike previous one-shot pruning methods, AST incrementally transforms dense models into sparse ones by applying decay to masked weights while allowing the model to adaptively select masks throughout the training process. Furthermore, we observe that using distillation with a dense model as the teacher can prevent the sparse model from falling into local optima and accelerate convergence. In addition, we incorporate extra well-initialized parameters to further enhance model performance with minimal increase in memory footprint. AST can significantly enhance model performance, approaching the level of dense models. When applied to the LLaMA2-7B model, AST reduces the zero-shot accuracy gap between dense and semi-structured sparse models to 1.12% across multiple zero-shot tasks, utilizing less than 0.4% of the pretraining tokens. Our work demonstrates the feasibility of deploying semi-structured sparse large language models and introduces a novel method for achieving highly compressed models when combined with existing quantization techniques.
1 Introduction
Transformer-based Large Language Models (LLMs) are equipped to handle complex tasks (Devlin et al. 2018; Brown et al. 2020; Achiam et al. 2023) and exhibit emergent abilities (Wei et al. 2022) due to their expanding parameter count. However, this expansion poses challenges to the deployment of current state-of-the-art models, as inference speed is constrained by increasing memory latency, leading to significant computational and memory requirements. This has spurred a series of efforts to develop effective model compression techniques to reduce memory footprint and lower deployment barriers.
Model pruning (Frantar and Alistarh 2023; Han, Mao, and Dally 2016; Sun et al. 2023) is an effective technique for compressing LLMs by setting a proportion of weights to zero, thereby adhering to a specific sparsity pattern. Recently, N:M sparsity has emerged as a type of semi-structured sparsity that offers an optimal balance between precision and hardware efficiency. Specifically, N:M sparsity retains only N nonzero elements out of every group of M elements. This sparsity pattern can accelerate both matrix multiplication and memory access, potentially enhancing the performance of both pre-filling and decoding processes on off-the-shelf GPUs.
Despite the promise of N:M pruning, current state-of-the-art methods for LLMs, such as SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023), employ a post-training approach that determines the sparsity pattern layer by layer without requiring backpropagation. Although this approach improves efficiency, it often leads to significant performance degradation, particularly in knowledge-intensive tasks (Nowak et al. 2024), raising concerns about the feasibility of pruning LLMs. Additionally, while retraining pruned models has been successful in the pre-LLM era (Wang, Wohlwend, and Lei 2019; Lee, Ajanthan, and Torr 2019; Evci et al. 2020), its application to models with billions of parameters remains underexplored.
In spite of its potential, retraining sparse LLMs presents several unique challenges: (1) Retraining is computationally expensive, particularly when compared to training-free pruning methods, necessitating techniques that ensure rapid convergence. (2) The absence of original pretraining datasets often forces reliance on weaker open-source datasets for retraining, which may not fully capture the knowledge encoded in the original model. (3) The model must conform to the strict requirements of semi-structured sparsity, which adds additional constraints to the retraining process. (4) The pruning of LLMs can result in the loss of crucial language understanding and reasoning abilities, which are considerably more difficult to recover than metrics like perplexity.
To overcome these challenges, we propose a novel and lightweight training method, Adaptive Sparse Trainer (AST), designed to produce high-performance sparse LLMs. For the first time, we demonstrate that semi-structured sparse LLMs can achieve competitive performance not only in terms of perplexity but also in more demanding knowledge-intensive tasks, making them viable for practical deployment. AST integrates the transition from dense to sparse models directly into the training process, gradually decaying unimportant weights to zero through a carefully designed decay mechanism. This approach allows for the revival of pruned weights during training, enabling a smooth and dynamic adjustment of the model’s connectivity pattern while adhering to the N:M sparsity structure. Furthermore, AST applies knowledge distillation using a dense model as the teacher, which can significantly speed up model convergence and prevent sparse models from getting stuck in local optima. It also helps the sparse model retain the valuable world knowledge and performance characteristics of the original dense model, thereby enhancing its generalization ability and compensating for the use of weaker training datasets. To further enhance performance, we introduce the use of a small number of carefully initialized parameters. By combining the RedPajama dataset with dynamic batching (Xia et al. 2024), AST effectively maximizes the potential of sparse models. When applied to the LLaMA2-7B model, AST achieves a 2:4 sparse configuration with minimal performance loss, demonstrating that compressed models can still perform effectively in practice. Our model further benefits from AWQ quantization, achieving competitive compression rates with state-of-the-art performance.
The key contributions of this work are as follows:
  • We propose Adaptive Sparse Trainer (AST), a novel semi-structured pruning framework designed to compress large language models efficiently while maintaining high performance.
  • AST introduces a gradual decay scheduler (Annealing SR-STE) combined with knowledge distillation to accelerate model convergence and improve performance on complex tasks.
  • We introduce SLoRB, a technique that boosts model performance by adding a small number of well-initialized parameters.
  • Applied to the LLaMA2-7B model, our method achieves a 2:4 sparse configuration with only a 0.6 increase in perplexity and a 1.2% accuracy drop in zero-shot tasks, using less than 0.4% of the original pretraining tokens.
2 Related Work
Network Pruning
Pruning is a crucial technique aimed at reducing model size and computational costs. It originated from methods like Optimal Brain Damage (LeCun, Denker, and Solla 1990) and Optimal Brain Surgeon (Hassibi, Stork, and Wolff 1993). Based on sparsity patterns, pruning methods can be broadly categorized into unstructured, structured, and semi-structured pruning. Unstructured pruning removes individual weights (Han, Mao, and Dally 2016; Paul et al. 2023), which can maintain performance even with high sparsity. However due to its random pattern, unstructured models are difficult to accelerate. Structured pruning (Liu et al. 2017; Molchanov et al. 2019; Nova, Dai, and Schuurmans 2023; Shen et al. 2022), on the other hand, removes entire neurons, filters, or attention heads, resulting in models that are easier to accelerate on standard hardware but often suffer from severe performance loss. Semi-structured sparsity (e.g., N:M sparsity) (Hubara et al. 2021) has been applied as a trade-off between performance and achieving actual speedup. Recently a series of works (Frantar and Alistarh 2023; Sun et al. 2023; Zhang et al. 2024a, b) have made progress in pruning LLMs with billions of parameters. However, these pruned models from training-free methods still fall short in complex zero-shot performance.
Retraining Pruned Models
Another line of research (Han et al. 2015; Singh and Alistarh 2020; Renda, Frankle, and Carbin 2020; Zhou et al. 2023) has focused on retraining pruned models to enhance performance. While retraining has been shown to be effective for smaller models and simpler tasks (Kurtic and Alistarh 2022; Zhu and Gupta 2018), it often involves additional training steps (Frankle and Carbin 2019) or introduces extra parameters (Shi et al. 2023) for pruning, which limits its application to large-scale models due to the substantial computational resources required. Other retraining methods focus on unstructured sparsity, which may struggle when given stricter sparsity patterns (Lee, Ajanthan, and Torr 2019; Evci et al. 2020). Recently, Sheared LLaMA (Xia et al. 2024) employed a two-stage structured pruning process to train models that outperform others of similar size. However, they employ a computationally intensive pruning stage unsuitable for more fine-grained sparsity. Our work proposes a lightweight pruning process that enables the retraining of semi-structured sparse large language models with minimal training costs.
Combining Pruning with Quantization
Pruned models can be further compressed using quantization. Earlier methods like Deep Compression (Han, Mao, and Dally 2016) combined pruning and quantization to significantly reduce the size of neural networks. More recent work (Frantar and Alistarh 2023; Sun et al. 2023; Guo et al. 2024) has combined sparsity with quantization in large language models. In our work, we report results using AWQ quantization (Lin et al. 2024) with our semi-structured sparse model.
3 Methods
Figure 1: (Left) In the naive training baseline, the mask remains constant during training, which can result in suboptimal performance. (Right) Adaptive sparse training enables both mask and weight learning through a scheduled decay term. AST also utilizes distillation and SLoRB parameters to speed up convergence and improve performance.
We begin by revisiting previous approaches to training sparse models, before introducing our method, the Adaptive Sparse Trainer (AST), which is described in detail. The full training pipeline is illustrated in Figure 1.
3.1 Sparsity Training
Pruning model weights is equivalent to multiplying them by an element-wise mask. For example, consider the matrix multiplication in a linear layer:
$Z=XW^{\top},\quad Z\in\mathbb{R}^{N\times D},\quad X\in\mathbb{R}^{N\times C},\quad W\in\mathbb{R}^{D\times C},$
where $X$, $W$, and $Z$ represent the model input, weight matrix, and output activation, respectively. The pruned weight matrix can be expressed as:
$\tilde{W}=m(W)\odot W,\quad m(W)\in\{0,1\}^{D\times C},$
where $m(\cdot)$ is a mapping that selects a mask based on the weight matrix $W$. In this work, we focus on N:M sparsity (Hubara et al. 2021), where the sparsity constraint ensures that there are $N$ nonzero elements in every $M$ consecutive weights connected to the same output.
However, when implementing the backward pass for a sparse model with automatic differentiation, the gradient cannot flow through the discrete mask $m(W)$. Previous work addresses this issue by using the straight-through estimator (STE) (Bengio, Léonard, and Courville 2013), which allows the gradient to pass through to the weights $W_{t}$ by ignoring the masking operation in the backward pass. This updates the parameters using the gradient with respect to the masked weight $\tilde{W_{t}}$, where $W_{t}$ is the parameter at iteration $t$:
$W_{t+1}\leftarrow W_{t}-\gamma_{t}g(\tilde{W_{t}}),$ (1)
here $g(\tilde{W_{t}})$ represents the gradient with respect to $\tilde{W_{t}}$ and $\gamma_{t}$ indicates the learning rate at iteration $t$.
One common approach for retraining LLMs with specific sparsity patterns, utilizing the Straight-Through Estimator (STE) (Sun et al. 2023), involves freezing the sparsity mask at the start of training and updating only the unmasked weights using the pretraining loss function, as depicted in Figure 1. However, we identified two significant limitations with this method. First, it discards the valuable information embedded in the pruned weights and prevents a smooth transition from a dense to a sparse model, resulting in slower convergence and suboptimal performance. Second, relying solely on the language modeling loss during retraining increases the risk of the sparse model becoming trapped in a local optimum, yielding poor outcomes.
[…]
3.2 Adaptive Mask Learning with Annealing SR-STE
As highlighted in previous work (Frankle and Carbin 2019; Evci et al. 2020), the connectivity pattern is as crucial as the weights themselves in determining the performance of sparse models. Therefore, we enable the model to adaptively learn the optimal mask during training instead of fixing it permanently. Specifically, we recalculate $m(W_{t})$ based on the magnitude criteria every $\Delta t$ iterations and add a decay term to the masked weights. The intuition behind this approach is straightforward: important weights typically receive strong gradients of the same sign during training, causing their magnitude to grow despite the decay, and they will eventually be revived. In contrast, less important weights tend to receive weaker, mixed gradient signals, leading them to decay to zero and remain masked under our settings.
However, selecting the appropriate decay strength is challenging. If the decay signal is too strong, masked weights will struggle to regenerate, leading to an almost fixed mask. On the other hand, if the decay signal is too weak, the model will fail to filter out important weights, resulting in heavy oscillation that hinders model convergence. To address this dilemma between stability and exploration, we propose adding a moving decay schedule, with weaker decay strength at the beginning to encourage the model to explore various connectivity patterns and a stronger decay signal toward the end to facilitate model convergence.
Formally, we update the model weights with the following equation:
$\lambda_{W}(t)=\begin{cases}\alpha t,&\text{if }0\leq t\leq T_{0},\\ \alpha T_{0},&\text{if }T_{0}\leq t\leq T_{1},\end{cases}$ (2)
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\left(g(\tilde{W_{t}})+\lambda_{W}(t)\left(\overline{m(W_{t})}\odot W_{t}\right)\right),$ (3)
where $\lambda_{W}(t)$ denotes the decay factor, $\overline{m(W_{t})}$ denotes the mask for pruned weights at iteration $t$, $T_{1}$ denotes the total number of training batches, and $T_{0}$ represents the batch at which the decay factor stops increasing. Compared with STE in Equation 1, masked weights receive a decay signal proportional to the decay factor $\lambda_{W}(t)$, which starts small at the beginning and increases later on.
This method, termed Annealing SR-STE (ARS-STE), is adapted from the standard SR-STE (Zhou et al. 2021). While their work focuses on pretraining settings and is primarily used to prevent mask oscillation, we apply it in the retraining context to filter out important weights. Applying ARS-STE has been shown to improve model performance compared to naive SR-STE. We also provide a detailed analysis of mask oscillation behavior during training in section 6.7.
We also experimented with different one-shot pruning criteria that incorporate activation information (Frantar and Alistarh 2023; Zhang et al. 2024a) for mask selection. However, we observed that the classic magnitude criterion consistently outperforms these methods during training, both in terms of computational efficiency and performance. Activation-based methods involve higher computational costs compared to magnitude-based masks. Moreover, their reliability diminishes as fluctuations in activation levels due to weight updates undermine their effectiveness, particularly in the later stages of training.
Given that training LLMs typically involves the use of the AdamW optimizer, we apply decay directly to the weights, as demonstrated in recent work (Hu et al. 2024). The detailed expressions and explanations are also included in Section 6.6.
3.3 Alleviating the Retraining Dilemma through Knowledge Distillation
A key distinction between pretraining and retraining pruned models lies in how their parameters are initialized. In the pretraining setting, parameters are typically randomly sampled, whereas in retraining, they are inherited from a well-trained model. Consequently, pruned models retain a portion of the patterns acquired during pretraining.
We discovered that this feature tends to trap pruned models in local optima. Specifically, while retraining can achieve much faster initial convergence, it often fails to reach an optimal state later on; we refer to this phenomenon as the Retraining Dilemma. As illustrated in Figure 2, although using cross-entropy loss during retraining initially leads to a quicker reduction in training loss compared to pretraining, the test set perplexity remains unstable and elevated. We attribute this to the fact that, unlike randomly initialized models, the pruned model is more prone to overfitting on the limited data available at the start due to its prior knowledge. This overfitting hampers the model’s ability to learn global features and achieve optimal convergence. Notably, this issue persists even when using a significantly smaller learning rate than in pretraining, suggesting that simply reducing the learning rate is insufficient to resolve the problem.
Figure 2: (Upper) The wikitext perplexity curve for retraining GPT2 with and without knowledge distillation. (Lower) The training loss curve for pretraining and retraining.
As a solution, we found that applying KL-divergence loss (Kullback and Leibler 1951) can address this issue. Unlike the language modeling loss, KL-divergence loss measures the difference between two probability distributions, offering richer signals that help mitigate overfitting in the early stages of training. Consequently, we employ the following loss function:
$\mathcal{L}_{logit}=D_{KL}(p_{\theta_{t}}||p_{\theta_{s}})$ $=\frac{1}{B\times S}\sum_{i=1}^{B\times\text{seq}}p_{\theta_{t}}(x_{i})\log\frac{p_{\theta_{t}}(x_{i})}{p_{\theta_{s}}(x_{i})},$
$\mathcal{L}=$ $\alpha\mathcal{L}_{logit}+(1-\alpha)\mathcal{L}_{task},$ (4)
where $\mathcal{L}_{task}$ is the cross-entropy loss, $B$ is the batch size, $S$ is the sequence length, and $p_{\theta_{t}}$ and $p_{\theta_{s}}$ are the probability distributions of the teacher and student models, respectively.
We observe that as Retraining Dilemma is more pronounced in smaller models; therefore, we typically apply a larger $\alpha$ for these models. We also experimented with various distillation methods that utilize intermediate information for distillation. However, we found that adding constraints to intermediate outputs hinders the model’s generalization ability and leads to undesirable outcomes.
3.4 Sparse Low-Rank Boosting
The pruned model with 2:4 sparsity, which retains only half of the parameters under strict structural constraints, experiences a reduction in expressive capacity. To mitigate this, we incorporate a LoRA (Low-Rank Adaptation) (Hu et al. 2022) shape parameter that is trained alongside the original parameters, helping to bridge the performance gap between dense and sparse models with minimal memory overhead. Unlike traditional LoRA fine-tuning, where the original model parameters are frozen and only the adapters are trained, our approach trains both the sparse parameters and the adapter weights simultaneously. This approach recognizes that while the classic LoRA method freezes the original model to prevent overfitting and catastrophic forgetting during downstream tasks, our method enhances generalization by training both the additional and original weights on a pretraining dataset, thereby enabling concurrent training.
Another notable aspect of our method is the initialization strategy. Classic LoRA typically employs random initialization techniques, such as Xavier initialization (Glorot and Bengio 2010) or Kaiming initialization (He et al. 2015). However, in our approach, we leverage the pruned weights as additional information to initialize the adapter weights, thereby accelerating the training process. Given that in 2:4 sparsity, each neuron can only access half of the inputs, we find that incorporating additional information to retain the weight’s first-order information can help preserve the model’s capacity. Specifically, for a weight matrix $W$ of size $N\times d$ and its corresponding mask matrix $M$, we select the rank $r$ as $\frac{d}{k}$, where $k$ can be 64, 32, 16, and so on. A smaller $k$ improves performance but also increases memory usage. We select a projection matrix $X$ of size $\frac{d}{k}\times d$ and a SLoRB weight matrix $S$ of size $N\times\frac{d}{k}$. The matrices $X$ and $S$ are defined as follows:
$x_{ij}=\begin{cases}1,&\text{if }i\cdot k\leq j\leq(i+1)\cdot k-1,\\ 0,&\text{otherwise},\end{cases}$ (5)
$S_{ij}=\frac{1}{k}\sum_{p=j\cdot k}^{(j+1)\cdot k-1}W_{ip}\cdot{\neg M_{ip}}.$ (6)
We define Group $G_{ij}$ as the elements from $j\cdot k$ to $(i+1)\cdot k-1$ in row $i$. As illustrated in Fig 3, each SLoRB weight $S_{ij}$ is broadcast within Group $G_{ij}$. By setting $S_{ij}$ as the mean of all pruned weights in Group $G_{ij}$, this design ensures that the total mean of groups $G_{ij}$ remains consistent after pruning. Our initialization method provides the model with a better starting point in training. Although different initialization methods may ultimately converge to similar outcomes given sufficient training steps, our method converges more rapidly, making it particularly advantageous when the computational budget is limited, as demonstrated in Section 6.4. We refer to this method as Sparse Low-Rank Boosting (SLoRB), which trades off memory overhead for enhanced performance. The complete pseudocode for our method is provided in Algorithm 1.
Figure 3: Visualization of SLoRB Initialization: Consider a weight matrix of size 3 by 8, with $k=4$. Weight $S_{ij}$ is broadcasted within group $G_{ij}$ after multiplication with the projection matrix.
Algorithm 1 Training Process for AST
Input: Weight of linear layers $\mathcal{W}^{(0)}$; Mask update frequency $\Delta t$; Total training iteration $T_{0}$; SLoRB parameter k; Decay increase iteration $T_{1}$;
1:  for $W^{(0)}\in\mathcal{W}^{(0)}$ do
2:     Initialize the mask $m(W^{(0)})$ based on magnitude;
3:     if Use SLoRB then
4:        Initialize SLoRB weight $S^{(0)}\in\mathcal{S}^{(0)}$ and $X^{(0)}\in\mathcal{X}^{(0)}$ by Equation 5 and 6;
5:     end if
6:  end for
7:  for t = 1,2,..$T_{0}$ do
[…]
9:        for $W^{(t)}\in\mathcal{W}^{(t)}$ do
10:           Update model mask mask $m(W^{(t)})$;
11:        end for
12:     end if
13:     Compute decay for term $\lambda_{W}(t)$ from Equation 2;
14:     Compute gradient $g(W^{(t)})$ by back-propagation on distillation loss $\mathcal{L}$ from Equation 4;
15:     Update weight $W^{(t)}$ by Equation 3;
16:     if Use SLoRB then
17:        Update $S^{(t)}$ and $X^{(t)}$ by gradient;
18:     end if
19:  end for
20:  return the pruned model.
4 Experiments
Table 1: Perplexity results on raw-WikiText2 on 2:4 sparsified language models. AST outperforms both training-free and other training-based methods with similar computational costs.
OPT GPT2
Method Training 125M 350M 1.3B 124M 350M 774M 1.5B
Dense - 27.76 22.00 14.62 29.95 21.72 19.43 17.40
SparseGPT 45.58 40.33 29.03 50.09 31.03 25.98 21.14
Wanda 60.91 50.16 23.92 115.64 63.71 49.97 30.44
Iterative Magnitude Pruning 38.37 30.29 23.94 40.08 29.86 24.31 20.83
Gradual Magnitude Pruning 31.51 25.98 16.78 33.48 24.83 22.01 18.96
AST-Naive(Ours) 30.22 24.65 15.85 32.23 23.65 21.29 18.33
AST-Boosted(Ours) 28.68 24.03 15.43 31.13 23.03 20.66 18.01
Table 2: Accuracy (%) of various open-sourced models on seven zero-shot tasks.
Models (Sparsity Pattern) Parameters BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
LLaMA2-7B (Dense) 6.7B 77.73 63.89 57.18 69.04 76.17 42.91 31.60 59.78
Sheared-LLaMA-2.7B (Dense) 2.7B 65.99 50.54 51.21 65.03 67.29 33.27 28.80 51.73
INCITE-Base-3B (Dense) 2.8B 67.40 52.34 47.91 62.98 67.68 31.74 27.60 51.09
Open-LLaMA-3B-v2 (Dense) 3.4B 65.89 55.69 52.12 63.59 68.34 34.32 26.00 52.13
LLaMA-7B (Sparse) 3.4B 73.21 61.34 54.86 66.18 70.24 35.68 31.80 56.19
LLaMA2-7B (Sparse) 3.4B 73.12 66.06 54.66 67.87 73.61 39.93 28.60 57.68
LLaMA2-7B (Sparse+SLoRB) 4.2B 75.04 66.06 55.24 68.48 74.91 41.11 29.40 58.62
4.1 Experiment Setup
Model Configurations. We report the performance of Adaptive Sparse Trainer (AST) across three different LLM model families: LLaMA2-7B (Touvron et al. 2023), OPT (Zhang et al. 2022), and GPT2 (Brown et al. 2020). We present results for two different sparsity patterns: AST-Naive, which adheres to a strict 2:4 sparsity pattern without additional weights, and AST-Boosted, which incorporates extra SLoRB weights. For AST-Boosted models, we selected $k=16$, introducing an additional 12.5% of parameters. Optimal hyperparameters were identified through a grid search, with the specific hyperparameters and training details provided in Section 6.3.
Data. For training smaller models like the OPT and GPT2 model families, we utilized the C4 dataset (Raffel et al. 2020). For the LLaMA2-7B model, we employed a more comprehensive dataset, RedPajama-v1111https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T, which encompasses data from seven domains: CommonCrawl, C4, GitHub, Wikipedia, Books, ArXiv, and StackExchange. Additionally, we leveraged the dynamic batching feature provided in the ShearedLLaMA (Xia et al. 2024) codebase.
Baseline. For the GPT2 and OPT models, we compare our methods with both training-free and training-based approaches. Among training-free methods, we include comparisons with SparseGPT (Frantar and Alistarh 2023) and Wanda (Sun et al. 2023). For training-based methods, given the lack of existing results for retraining generative language models with N:M sparsity, we implemented iterative magnitude pruning (Frankle and Carbin 2019) and gradual magnitude pruning (Kurtic and Alistarh 2022), both of which can be adapted to achieve the desired sparsity. To ensure a fair comparison in terms of computational cost, we report results for training-based methods that include distillation, as we have observed that incorporating distillation significantly enhances performance. Due to computational constraints, we report the results of training-based baselines only for the GPT2 and OPT models.
For the LLaMA2-7B model, we compare our approach against strong dense pre-trained LLMs with a similar number of non-zero parameters. Additionally, we include a comparison with the LLaMA-7B 2:4 sparsity model provided in the Wanda (Sun et al. 2023) paper.
Evaluation. We evaluated the WikiText-2 perplexity for all models and assessed zero-shot and few-shot performance on LLaMA2-7B using EleutherAI’s LM Harness (Gao et al. 2021). Our evaluation tasks included zero-shot ARC Easy (Clark et al. 2018), OpenBookQA (Mihaylov et al. 2018), WinoGrande (Sakaguchi et al. 2021), RTE from the GLUE benchmark (Wang et al. 2018), HellaSwag (Zellers et al. 2019), ARC Challenge (Clark et al. 2018), BoolQ (Clark et al. 2019), as well as more complex tasks like the MMLU benchmark (Hendrycks et al. 2021a), GSM8K (Cobbe et al. 2021), and MATH (Hendrycks et al. 2021b).
4.2 Main Results
Language Modeling In Table 1, we compare the perplexity results of our sparse models with those of the baselines. Our methods significantly outperform both training-free and training-based approaches across various models, achieving results close to those of the dense baseline. A plausible reason why training-based methods like IMP fail is that while a lottery ticket may exist for a randomly initialized model, this may not be the case for a well-trained model. As for the gradual baseline, applying a proportion of training steps at low sparsity may not provide the model with sufficient time to converge.
Zero-shot Result In Table 2, we present the accuracy results of seven zero-shot tasks for AST and baseline models. Our method performed best in most of the tasks. It should be noted that our sparse models consistently outperform smaller dense models with a similar parameter count, suggesting a promising approach for obtaining parameter-efficient models. Moreover, our method requires only a minimal amount of training tokens to converge. For example, when training the LLaMA2-7B model, we utilized only 7B tokens, which is less than 0.4% of those used in pretraining, making AST applicable to open-source models.
Our method can also be seamlessly adapted to current quantization techniques to achieve extremely compressed models with minimal performance drop. We provide results using AWQ quantization methods in section 6.8.
4.3 Additional Results for LLaMA2-7B
Recent findings (Nowak et al. 2024) have shown that previous pruning methods suffer significant performance degradation in knowledge-intensive tasks. To address this, we tested our LLaMA2-7B model on more complex tasks. As shown in Table 3, our model preserves most of the knowledge from the dense model and continues to perform well in knowledge-intensive tasks compared with previous pruning methods. This provides strong evidence that 2:4 pruned models possess considerable potential, contrary to the observations in previous studies.
Table 3: Results of perplexity and knowledge-intensive tasks for LLaMA2-7B models with 2:4 sparsity. The symbols $\downarrow$ and $\uparrow$ indicate that lower and higher values are better, respectively. (LoRA methods are finetuned with $r$=64)
LLaMA2-7B Dense Wanda AST-Naive AST-Boosted
Perplexity $\downarrow$ 5.12 11.02 5.82 5.69
MMLU (5-shot) $\uparrow$ 45.3 27.6 37.9 38.2
MATH (4-shot) $\uparrow$ 5.38 2.86 4.42 4.64
GMS8K (LoRA Finetuned) $\uparrow$ 40.3 32.1 35.6 36.2
4.4 Ablation Study
Training Sparse Models. Our AST-Naive method employed distillation, Annealing SR-STE, and adaptive masking during training. In Table 4, we present an ablation study to assess the effects of each of these components. For a fair comparison, we apply additional training steps to non-distillation methods. Specifically, we analyze the impact of training without distillation, using a naive SR-STE decay factor, and fixing the mask during training, each component in isolation. Additionally, we provide results for the naive training baseline in Figure 1 mentioned above. We demonstrate that all three approaches contribute to performance gains compared to naive training across various models.
Table 4: Ablation study of different methods in training sparse models.
GPT OPT
Method 124M 350M 774M 125M
Dense 29.95 21.72 19.43 27.76
Naive Training 40.34 29.79 28.65 39.46
No distillation 39.29 29.08 27.21 36.97
Static SR-STE 32.84 24.04 21.73 31.08
Fixed Mask 32.93 24.18 21.95 31.06
AST-Naive(Ours) 32.23 23.65 21.29 30.22
Distillation Functions We also conducted ablation studies on different distillation functions used in previous work, including TinyBERT (Jiao et al. 2020), MobileBERT (Sun et al. 2020), and Sparse-Finetuning (Kurtic et al. 2023), which use attention and hidden states for distillation. Additionally, we examined MiniLLM (Gu et al. 2024), which employs reverse KL divergence. We find that using intermediate information during the distillation of generative language models is detrimental, and that KL loss is sufficient for optimal performance. Reverse-KL yields performance similar to that of forward KL. Detailed descriptions of the distillation loss functions for each of these methods are provided in section 6.5.
Table 5: Ablation study of different distillation loss for training sparse models.
GPT OPT
Method 124M 350M 774M 125M
Dense 29.95 21.72 19.43 27.76
TinyBERT 42.75 33.35 30.86 39.39
MobileBERT 44.87 31.67 29.75 40.33
Sparse-Fintuning 41.19 29.42 26.19 38.96
MiniLLM 32.20 23.64 21.31 30.24
KL Loss(Ours) 32.23 23.65 21.29 30.22
4.5 Speedup
In Table LABEL:tab6, we report speedup results using PyTorch support for 2:4 sparse models [^2] https://pytorch.org/docs/2.1/sparse.html#sparse-semi-structured-tensors. Since the current API only accelerates computation without reducing memory consumption, we focus on reporting results for matrix multiplication and end-to-end inference on LLaMA2-7B. 2:4 sprase model can achieves up to a 1.91x speedup in matrix multiplication and up to 1.35x in end-to-end inference.
Table 6: Speedup results using PyTorch API. We fixed the size of the sparse matrix to be 4096 by 4096 and varied the input matrix size when testing matrix multiplication.
Matrix Multiplication
Block Size Sparse Dense Speedup
12288 3.80ms 1.99ms 1.91x
8096 2.54ms 1.33ms 1.91x
4096 1.28ms 0.68ms 1.88x
End-to-End
Seqlen Sparse Dense Speedup
2048 289.07ms 244.80ms 1.18x
1024 142.04ms 105.22ms 1.35x
512 73.84ms 55.68ms 1.35x
5 Conclusion
In this paper, we introduce the Adaptive Sparse Trainer (AST), a novel training pipeline designed for semi-structured sparse models. AST effectively narrows the precision gap between dense and sparse large language models (LLMs), both in terms of perplexity and performance on knowledge-intensive tasks, while maintaining minimal training costs. Our work demonstrates the feasibility of pruning LLMs without significant performance loss, showing that large semi-structured sparse models can outperform dense models of similar size when paired with appropriate hardware. Additionally, AST achieves substantial performance gains without compromising inference speed, highlighting its potential for the efficient deployment of sparse LLMs. While our findings contribute to the advancement of retraining pruned models with billions of parameters, we acknowledge that the use of a limited number of training tokens due to computational constraints is an area for future exploration. Extending this work to larger models or increasing the number of training tokens could provide valuable insights and further enhance the effectiveness of AST.
6 Appendix
6.1 Compressing Models with Different Semi-structured Sparsity Pattern
In previous sections, we only examined 2:4 patterns, but theoretically, any N:M pattern can enhance matrix multiplication with appropriate hardware support. With that said, patterns with looser constraints (i.e., a larger M) often perform better but are more challenging to compress effectively. We aim to identify the optimal trade-off and. To maintain fairness in comparison, we focus only on patterns that achieve a 50% sparsity ratio, such as 2:4, 4:8, etc.
It should be noted that deploying semi-structured sparse models can lead to potential memory overhead. Take the 2:4 sparsity pattern as an example: although only two non-zero elements out of every four weights need to be stored, additional positional information is required to correctly restore these compressed weights to the original matrix for calculations. Storing these indices will consume an additional proportion of memory. A naive method would be to store one bit as a mask for each element, which results in a 50% memory overhead with a 4-bit quantized model at 50% sparsity. However, a more efficient approach exists: since there are 2 non-zero elements, there are a total of (${4\choose 2}=6$ possible positions, hence only an additional 3 bits are sufficient.
We define the compression ratio as the percentage of memory latency of the compressed model compared to the original model. In our setting, we compress dense models with FP32 precision to a 4-bit model with $n$:$2n$ semi-structured sparsity; therefore, the compression ratio can be formulated as:
$C_{n}=\frac{n*4+\lceil\log_{2}{{2n\choose n}}\rceil(bit)}{2n*32(bit)}=\frac{1}{16}+\frac{\lceil\log_{2}{{2n\choose n}}\rceil}{64n}.$
Through mathematical formulation (provided in Section 6.2), we can provide an approximate upper bound for the compression ratio as well as a predicted actual compression ratio when extending to sparsity patterns with larger $N$ and $M$
$\log_{2}{2n\choose n}=\log_{2}\frac{(2n)!}{(n!)^{2}}\leq\log_{2}\frac{4^{n}}{\sqrt{\pi n}}=2n-\log_{2}\sqrt{\pi n}.$
If we remove the ceiling function for approximation:
$C_{n}\approx\frac{3}{32}-\frac{\log_{2}\sqrt{\pi n}}{64n}.$
Therefore, the compression ratio is approximately an increasing function with an upper bound of $C^{*}=\frac{3}{32}\approx 9.375\%$. In practice, to further enhance compression, we can employ Huffman Encoding, similar to what was done in Deep Compression (Han, Mao, and Dally 2016). We construct a Huffman Tree with evenly distributed frequencies, as empirical results support our claim. The actual compression ratios are shown in Table 7. We observe that as $n$ increases, performance improves; however, although the compression ratio increases, it remains below the upper bound.
Table 7: Comparison of perplexity and compression ratio of different sparsity pattern.
Sparsity Pattern 1:2 2:4 4:8 8:16 16:32 32:64 Unstructured 50%
Perplexity 32.56 31.13 30.73 30.37 30.34 30.32 30.18
Actual Compression Ratio 7.81% 8.33% 8.66% 8.93% 9.10% 9.22% -
6.2 Proof for Upper Bound of Combination Number
Proofs are sourced from Stack Exchange  [^3] https://math.stackexchange.com/questions/1442351/stirling-approximation-of-binom2nn. We have the follow equation:
$\frac{1}{4^{n}}{2n\choose n}=\frac{(2n-1)!!}{(2n)!!}=\prod_{k=1}^{n}(1-\frac{1}{2k}).$
We square both sides of the equation:
$(\frac{1}{4^{n}}{2n\choose n})^{2}=\frac{1}{4}\prod_{k=2}^{n}(1-\frac{1}{2k})^{2}=\frac{1}{4n}\prod_{k=1}^{n-1}(1-\frac{1}{(2k+1)^{2}})^{-1}.$
Using the Weierstrass product for the cosine function, we obtain:
$\prod_{k=1}^{\infty}(1-\frac{1}{(2k+1)^{2}})^{-1}=\frac{4}{\pi}.$
Hence, it follows that:
$(\frac{1}{4^{n}}{2n\choose n})^{2}=\frac{1}{\pi n}\prod_{k\geq n}^{n}(1-\frac{1}{(2k+1)^{2}})=\frac{1}{\pi n}\prod_{k\geq n}^{n}(1+\frac{1}{(2k+2)2k})^{-1}.$
Therefore we have
$\frac{1}{\sqrt{\pi n}}\geq\frac{1}{4^{n}}{2n\choose n}.$
6.3 Hyperparamters
We present the hyper-parameters and the number of tokens used to train our model in Table 8. The increase iteration number $T_{0}$ is selected to be one-fourth of the total training iteration $T_{1}$.
Table 8: Summary of hyperparameters and the number of tokens used for training.
OPT GPT2 LLAMA2
125M 350M 1.3B 124M 350M 774M 1.5B 7B
Learning Rate 1e-4 1e-4 2e-5 2e-4 1e-4 1e-4 6e-5 2e-5
Maximal Decay Factor 3e-4 1e-4 6e-5 1e-4 6e-5 6e-5 6e-5 6e-5
Training Steps 40k 40k 20k 40k 40k 40k 20k 15k
Total Flipped Ratio 11.1% 9.3% 4.4% 6.7% 3.6% 3.4% 4.2% 3.8%
Kl loss parameter $2/3$ $2/3$ $2/3$ $2/3$ $2/3$ $2/3$ $2/3$ $1/3$
Tokens Trained 10B 10B 5B 5B 5B 5B 2.5B 7.5B
6.4 SLoRB Initialization
We conducted an ablation study on various initialization methods for SLoRB. We report the results on GPT2 model under specified computation limits of 1B, 2.5B, and 5B tokens, as shown in Table 9. ’Mean’ refers to the initialization method discussed in the paper, while ’Zero’ indicates the initialization of both matrices from zero. Our initialization method assists in maintaining model performance at the outset, thereby accelerating convergence.
Table 9: Results of perplexity using different initialization methods with fixed training tokens.
Trained Token 0 1B 2.5B 5B
Mean 475.05 32.78 31.78 31.13
Xavier Uniform 720.74 33.22 31.82 31.15
Zero 720.74 33.43 32.93 32.46
6.5 Different Distillation Function
In this section, we formally present the loss functions applied in our ablation study, focusing on different distillation methods. We include results from previous works such as TinyBERT (Jiao et al. 2020), MobileBERT (Sun et al. 2020), Sparse-Finetuning (Kurtic et al. 2023), and MiniLLM (Gu et al. 2024).
TinyBERT
In TinyBERT, we utilize information from intermediate multi-head attentions and hidden states using the mean square error (MSE) loss. Specifically, the attention loss is defined as:
$\mathcal{L}_{attn}=\frac{1}{h}\sum_{i=1}^{h}\text{MSE}(A_{i}^{S},A_{i}^{T}),$ (7)
where $A_{i}^{S}$ and $A_{i}^{T}$ represent the layer-wise attention outputs of head $i$ from the student model and the teacher model, respectively.
The hidden state loss is defined as:
$\mathcal{L}_{hidn}=\text{MSE}(H^{S}W_{h},H^{T}),$ (8)
where $H^{S}$ and $H^{T}$ represent the layer-wise hidden states of the student model and teacher model, respectively. Here, $W_{h}$ is a learnable projection matrix. However, since the shapes of the hidden states are identical in our settings, we fix $W_{h}=I$.
TinyBERT also uses cross-entropy loss:
$\mathcal{L}_{pred}=\text{CE}(z^{T}/\tau,z^{S}/\tau),$ (9)
where $z^{T}$ and $z^{S}$ are the output logits of the teacher and student models, respectively. We do not utilize embedding distillation, as the embeddings are well-trained and not pruned.
The distillation function for each layer is defined as:
$\mathcal{L}_{layer}=\begin{cases}\mathcal{L}_{pred}&\text{if it is the last layer},\\ \mathcal{L}_{hidn}+\mathcal{L}_{attn}&\text{otherwise}.\end{cases}$ (10)
The final loss function is the sum of the layer-wise loss functions.
MobileBERT
MobileBERT is designed for distilling BERT by utilizing information from feature maps and attention. The feature map distillation in MobileBERT is identical to that used in TinyBERT, as shown in Equation 8. For attention knowledge transfer at layer $l$, the loss function is defined as:
$\mathcal{L}_{AT}^{l}=\frac{1}{AT}\sum_{t=1}^{T}\sum_{a=1}^{A}D_{KL}(a_{t,l,a}^{tr}\|a_{t,l,a}^{st}),$ (11)
where $A$ is the number of attention heads, $T$ is the sequence length, and $a_{t,l,a}$ represents the attention output for layer $l$, head $a$, and token $t$. The total distillation loss is the sum of the feature map loss and attention transfer.
Sparse-Finetuning
Sparse-Finetuning employs both SquareHead Loss and KL-divergence loss. It measures the intermediate representation using mean squared error (MSE):
$\mathcal{L}_{feat}=\frac{\text{MSE}(f_{t}^{l},f_{s}^{l})}{\text{MSE}(f_{t}^{l},0)},$ (12)
where $f_{t}^{l}$ and $f_{s}^{l}$ are the hidden states of the $l$-th layer at step $t$ for the teacher and student models, respectively.
KL-divergence loss (Kullback and Leibler 1951) is used to measure the difference in output probability distributions between the student and teacher models:
$\mathcal{L}_{logit}=D_{KL}(p_{\theta_{t}}\|p_{\theta_{s}}),$ (13)
where $p_{\theta_{t}}$ and $p_{\theta_{s}}$ are the probability distributions of the teacher and student models, respectively.
The total distillation loss is the weighted sum:
$\mathcal{L}=\alpha_{1}\mathcal{L}_{logit}+\alpha_{2}\mathcal{L}_{feat}.$ (14)
MiniLLM
MiniLLM uses reverse KL-divergence instead of forward KL. Specifically, the loss function is defined as:
$\mathcal{L}_{logit}=D_{KL}(p_{\theta_{s}}\|p_{\theta_{t}}),$ (15)
where $p_{\theta_{t}}$ and $p_{\theta_{s}}$ are the probability distributions of the teacher and student models, respectively.
6.6 Adding Decay for Adam Optimizer
Recent work (Hu et al. 2024) has discovered that using SR-STE with momentum-based optimizers can diminish the effectiveness of the decay term, leading to suboptimal results. Previously, the decay term was added to the gradient as follows:
$g_{t}\leftarrow\tilde{g_{t}}+\lambda_{W}(\overline{m(W_{t})}\odot W_{t}),$ (16)
$W_{t+1}\leftarrow W_{t}-\frac{\gamma(u_{t}\beta_{1}+(1-\beta_{1})g_{t})}{(1-\beta_{1}^{t})(\sqrt{\hat{v_{t}}}+\epsilon)},$ (17)
where $\tilde{g_{t}}$ is the gradient of $\tilde{W_{t}}$, and $u_{t}$ and $v_{t}$ are the first-order and second-order momenta, respectively.
Now, we add the decay term directly to the weight. First, we calculate the gradient $g(\tilde{W_{t}})$ using the Adam optimizer, and then update the weight using:
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\left(g(\tilde{W_{t}})+\lambda_{W}(\overline{m(W_{t})}\odot W_{t})\right).$ (18)
6.7 Mask Oscillation Behavior During Training
Several metrics have been introduced to measure the stability of masks during training, such as flip rate (Hu et al. 2024) and SAD (Zhou et al. 2021). In this paper, we primarily focus on measuring the flip rate and initial flip rate, which quantify the mask changes between the current and previous steps, as well as between the current and initial steps, respectively.
$r_{t}=||m(W_{t})-m(W_{t-1})||_{1}/D$ (19)
$i_{t}=||m(W_{t})-m(W_{0})||_{1}/D$ (20)
where $D$ represents the total number of parameters in the linear layer. The flip rate reflects the stability of the model masks, while the initial flip rate indicates the overall extent of weight removal and revival.
Table 10: Wikitext perplexity for quantized sparse models.
LLAMA2 OPT GPT2
Method
Theoretical
Compression Rate
7B 125M 1.3B 124M 350M 774M 1.5B
Dense 1.0x 5.12 27.76 14.62 29.95 21.72 19.43 17.40
AWQ-4bit 0.125x 5.68 29.12 14.95 31.93 22.66 19.89 17.80
AWQ-2bit 0.0675x 2.2e5 216.11 53.42 751.15 453.45 70.36 46.17
AST-Naive-8bit
0.125x 6.37 30.26 15.86 32.36 23.66 21.29 18.34
AST-Naive-4bit
0.0675x 6.48 31.28 16.05 34.00 24.39 21.72 18.60
AST-Boosted-4bit
0.078x 6.25 29.88 15.63 32.76 23.87 21.40 18.31
We compare the flip rates and initial flip rates of static SR-STE and Annealing SR-STE. In our experiments, the mask is updated, and the flip rate and initial flip rate are calculated every 10 batches during training, as more frequent recalculations do not improve accuracy and only increase computational overhead. Figure 4 illustrates the flip rates and initial flip rates during the retraining of the GPT2 model. Unlike traditional static decay factors, Annealing SR-STE modifies a higher percentage of the mask at the beginning, allowing the model to explore various mask patterns. Additionally, Annealing SR-STE enables a more stable mask towards the end of training, promoting better convergence. As a result, Annealing SR-STE supports a higher rate of mask changes (e.g., initial flip rate) while maintaining overall stability.
Figure 4: (Upper) Flip rate for static and Annealing SR-STE during the training of the GPT2 model. (Lower) Initial flip rate for static and Annealing SR-STE during the training of the GPT2 model.
6.8 Model Compression Using AWQ
In this section, we present the perplexity results for quantized models using AWQ. As shown in Table 10, our model maintains performance even under extreme compression. For the AST-Boosted model, we also quantized the additional SLoRB parameters.